# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ 4.5. Economic-driven engagement
Pak’s projects deliberately intertwine economic mechanisms with artistic interaction, turning market dynamics into part of the artistic experience. In many cases, the engagement of the audience is driven by financial gameplay—bidding wars, token burns, price competition—such that market participation becomes a form of performance. The Sotheby’s sale The Fungible is a paradigmatic example: it deployed a complex auction structure with open editions, surprise price hikes, and contests for unique rewards, effectively a “gamified” economic performance. This not only generated over \$20 million in sales, but also created a frenzy in which “more than 3,000 individuals decided to join,” drawn in by the competitive game itself. The mechanics of scarcity and reward (e.g. top buyers receiving special NFTs) were designed to incentivize maximal participation. Here the market behavior was the art: once “the game and competition incentives disappear,” what remains are tokens whose value was defined by that performative context . Pak thus highlights, even cynically, the extent to which economic structures confer meaning in crypto art.
@@
In Pak’s oeuvre, the act of buying or transacting becomes a critical ritual and medium. The Title (2020) consisted of identical images sold at different prices, making price itself the distinguishing feature of the artwork. One edition titled “Unsold” was listed for \$1 million and intentionally left unsold – a conceptual gesture underlining how value in art can be a product of hype and perception. This recalls Yves Klein’s 1957 exhibition of identical blue canvases sold at varying prices; as Klein observed, “the price…legitimately changes the experience of the work” despite no physical difference. Similarly, Pak has argued that an NFT is like a currency – judging it by the image alone is as foolish as picking a banknote. By merging art and economics, Pak invites collectors to reflect on their own motivations – are they driven by aesthetic appreciation or speculative gain? The economy around the artwork becomes integral to its narrative, effectively turning collectors into players in a financial drama. In Pak’s protocol art, market engagement is not a byproduct but a core feature that actively shapes the work itself.
@@ 4.6. Poetic message embedding in interaction rituals
Beyond the technical and economic layers, Pak’s works often carry metaphorical or poetic messages that are realized through the audience’s interactive rituals. In other words, the process a participant must follow is itself laden with symbolic meaning. Critic A.V. Marraccini notes that Pak treats the NFT medium as “a way to engage audiences in participatory conceptual art” – the concepts emerge through what the viewers or users are made to do. For instance, burn.art turns the destructive act of burning NFTs into a creative ritual. The platform’s mantra “Burn art to get ashes to get art to burn art” is a circular poem in itself, invoking themes of death and rebirth. Participants enact a cycle of sacrifice and renewal: by destroying one token, they generate another (the \$ASH token), which can then be spent on new art or even burned again. This cyclical interaction is a performative allegory about value and transformation – literally illustrating creation-through-destruction in a way that words alone could not.
Pak’s collaborations also embed messages in their structure. Censored (2022), created with Julian Assange, is explicitly about free expression and censorship. The project had two parts – one a dynamic single edition (Clock) counting Assange’s days behind bars, and the other an open invitation for people to speak. By telling users that Censored is “about you” and allowing them to tokenize any message they wish, Pak made the audience’s personal expression the core of the piece. The very name Censored reminds participants that their recorded messages might be obscured or blacked out, forcing reflection on the power of speech under observation. In Lost Poets, Pak evokes Jorge Luis Borges’s Library of Babel: an infinite library of all possible texts. The project “enlists participants to imagine” the lost contents of that library by naming their Poet NFTs and eventually inscribing words onto them. The ritual of naming and writing thus becomes a meditation on how meaning is generated – each user’s creative act is a fragment of a larger, hidden poem. Pak uses these game-like interactions not just as gimmicks but to convey ideas: the audience realizes the “poetic” concept by performing it. This approach recalls aspects of Fluxus or conceptual art, where simple actions (publishing a message, paying an exorbitant sum) become symbolic acts. The blockchain in Pak’s work provides a global, participatory stage for such rituals, ensuring that each interaction – each burn, each tokenized word – is both an artwork and a narrative gesture in his conceptual universe.
@@ 4.7. Interoperability enabling composability for emergence
Finally, Pak’s protocol-based approach leverages the interoperability of blockchain tokens to compose complex emergent ecosystems that transcend any single artwork. In blockchain terms, assets and smart contracts are composable: they can interact, integrate, or build upon each other permissionlessly. Pak exploits this composability to create an interlinked network of projects. The clearest example is the Ash token ecosystem. After The Fungible, Pak introduced \$ASH as a cryptocurrency earned by burning NFTs on burn.art. This token was not confined to one piece, but became a connective thread between works – a meta-artwork in its own right. Owners of ASH were rewarded in subsequent projects: before launching Lost Poets, Pak took a snapshot and airdropped 7,586 Pages to collectors holding at least 25 \$ASH, thus bridging the burn.art economy with the new literary NFT game. Conversely, once Lost Poets concluded, those Poet NFTs could be burned for ASH, feeding value back into the burn.art system. In this way, Pak created a feedback loop: one artwork’s output becomes another’s input. The emergent whole is an ecosystem of NFTs and tokens co-evolving across his oeuvre.
@@ 5. Discussion
5.1. Answering RQ: Characteristics of Protocol Arts
@@
This study’s exploration of Pak’s protocol artworks reveals seven interlinked characteristics that distinguish protocol art as a practice. (1) These works are fundamentally system-centric rather than object-centric. Their artistic meaning emerges from dynamic interconnectedness via protocols instead of static artifacts. (2) As Galloway claimed ”protocol is how control exists after decentralization” (Galloway, 2001), protocol arts employ autonomous governance for open-ended control, often via self-executing code or decentralized mechanisms, so that the artwork’s evolution is not predetermined by a single author but can continue to unfold indefinitely (3) This leads to distributed agency and communal authorship, as creative influence is spread across many participants and technical actants rather than residing in one creator, aligning with the idea that agency (and thus authorship) is shared among human and non-human actors in a network. (4) Moreover, these artworks exhibit temporal dynamism and lifecycle aesthetics: they develop through time, with phases of growth or change designed into the experience, so that process and feedback become central to their aesthetic. (5) Engagement with such works is often economic-driven, integrating incentive structures (for example, token rewards or market dynamics) that actively shape how audiences interact – a phenomenon noted by art scholars observing the immediate economic incentives at play in crypto art. (6) At the same time, there is a deliberate embedding of poetic messages in interaction rituals: the very actions participants perform (trading, voting, combining elements, etc.) are imbued with symbolic meaning, turning user interactions into ritualized performances that carry the artwork’s conceptual message. (7) Finally, protocol artworks leverage interoperability enabling composability for emergence, meaning they are designed to plug into larger ecosystems of code and community; like modular components, they can be reconfigured or linked with other protocols, allowing unexpected behaviors and creative outcomes to emerge from these combinations. Collectively, these seven characteristics illustrate how protocol art shifts the locus of art-making from singular objects to evolving systems of interaction, value and meaning.
@@
By situating these findings in broader intellectual contexts, we can appreciate protocol art as part of an interdisciplinary conceptual lineage. The system-centric focus strongly resonates with complex systems theory and the “systems aesthetics” tradition in art: as Jack Burnham observed, modern art increasingly “does not reside in material entities, but in relations between people and between people and the components of their environment”. In protocol art, the artwork is essentially a complex adaptive system (Buckley, 2017) – with inputs, outputs, and feedback loops – which accords with theoretical models where adaptive behavior and emergent order arise from networked interactions over time. Its open-ended governance and temporal evolution connect to the field of artificial life, which seeks systems that continually produce novelty rather than reaching a fixed end state. This open-endedness (Stanley et al., 2017), coupled with deliberate feedback mechanisms, echoes the temporal dynamics of complex adaptive systems studied in complexity science, reinforcing how unpredictable, ongoing change becomes an aesthetic virtue. Meanwhile, the ethos of distributed agency and communal authorship in these works finds precedent in media theory and anthropology: for instance, actor-network theory contends that creative agency is distributed across a web of human and non-human actants, not concentrated in an autonomous individual. Similarly, in interactive and participatory art literatures, authorship is often viewed as an emergent, collective process – the community of participants effectively co-creates and continually reshapes the piece (Filippi, 2025). The inclusion of economic-driven engagement situates protocol art in dialogue with social computing and crypto-economic systems: just as Web3 platforms rely on token-based incentives to drive user participation and loyalty, protocol artworks harness economic game dynamics as artistic material, blurring the line between aesthetic experience and market behavior. This integration of economic and social incentives also invokes evolutionary game theory within complex systems, where value-based choices influence the system’s trajectory. Furthermore, the poetic messaging through interaction rituals can be interpreted via performance studies and ritual theory in media. Even routine or rule-bound interactions carry expressive and symbolic weight – indeed, ritualized acts are “anything but purposeless” and can “constitute social reality” through shared symbolic action. In protocol art, the choreography of user interactions (such as collective decision-making or repetitive transactions) functions as a form of narrative or commentary, akin to a ceremonial performance that conveys meaning beyond its practical function. Finally, the principle of interoperability and composability aligns with discussions in software and platform studies about modular design and emergent behavior. In blockchain-based art, for example, smart contracts are conceived as interoperable building blocks, combinable like Lego pieces to create novel structures (Sun et al., 2024). This composability not only reflects a technical design philosophy but also fosters creative emergence: new artistic forms and communities can arise by linking protocols together, much as complex behaviors emerge when simple units interact in networked ecologies. Linking these characteristics to established frameworks thus shows that protocol art is not an isolated novelty but converges with long-running threads in complexity science, new media art, and socio-technical theory, reframing them in a unique artistic context.
@@ 5.2.1. Beyond Conceptual Art
Joseph Kosuth’s One and Three Chairs (1965) established that an idea—framed through language and documentation—can eclipse the physical object. Pak embraces the primacy of ideas but transports them into a rule-executing substrate. The Title (2021), whose only variables are price and on-chain title, echoes Kosuth’s linguistic gambit, yet its value is continuously arbitrated by decentralized markets rather than by academic discourse or institutional framing. Because the smart contract is immutable, its “statement” cannot be re-contextualised by curators; it is enacted in perpetuity by the network. Thus Pak’s work is not merely an illustration of a concept but a permanently operational logic that strangers must literally transact with. The artwork lives as executable code—a self-verifying proposition—rather than a documented proposition that awaits institutional re-presentation.
@@ 5.2.3. Beyond Participatory / Crowdsourcing Art
Relational-aesthetic projects of the 1990s (Bourriaud) invited viewers to co-produce meaning through convivial encounters, but those encounters were typically ephemeral and locally bounded. Pak’s participation is ledger-bound. In Lost Poets (2021–22), each name or verse a collector inscribes becomes an indelible block in a public archive—a form of distributed authorship impossible in earlier participatory works whose traces remained in wall labels or catalogs. Crucially, participation is also economically weighted: burning a Page to earn words entails an explicit cost, so creative decisions carry financial risk. This introduces game-theoretic depth absent from classic crowdsourcing art (e.g., MTurk drawings or open Wikis), aligning Pak’s practice with emerging literatures on cryptoeconomic design rather than with social-practice documentation alone.
@@ 5.2.4. Beyond Interactive Art
Interactive art from Myron Krueger’s Videoplace to Rafael Lozano-Hemmer’s public-space works invites real-time feedback loops, but those loops terminate once the exhibition ends. Pak designs durable interaction. When a collector merges two Merge tokens or burns an NFT for \$ASH, the blockchain records a state change that affects every future viewer and holder. Interaction ceases to be a momentary spectacle and becomes infrastructure—a permanent mutation in the artwork’s ontology. Moreover, interaction is asynchronous and scalable: a wallet signing a transaction at 3 a.m. in Seoul alters the same global state later queried by a viewer in São Paulo, creating what media theorist Yuk Hui calls a “cosmotechnical” network of relations that far outstrips gallery-based interactivity.
@@ 5.2.5. Beyond Performance Art
Performance art often foregrounds the artist’s body in time—think of Abramović’s Rhythm 0. Pak removes the corporeal performer and installs a protocol that performs indefinitely. Hate (2021) is a one-day happening in the sense that Pak gifts and recalls untransferable tokens, yet the contract he published remains a template other artists can fork, extending the performance beyond the artist’s presence. Likewise, Censored (2022) transforms a two-day mint into an enduring political performance: Clock keeps ticking so long as Assange is imprisoned, delegating the durational aspect to automated metadata calls and DAO stewardship. Thus the locus of endurance shifts from flesh to code, echoing cybernetic performance theories but grounded in irreversible cryptographic commitments.
5.3. Limitation and Future Works
Our case study of Pak’s nine landmark protocol artworks demonstrates the unique characteristics how smart-contract logic can serve as an artistic medium, but it also exposes the field’s present fragmentation. By limiting our scope to a single creator, we reveal both the promise and the incompleteness of today’s scholarship: there is still no shared vocabulary, taxonomy, or methodological toolkit robust enough to account for the full diversity of protocol-driven practice. We therefore call for a concerted, interdisciplinary effort to systemize knowledge around protocol art. Concretely, future work should

### v2

@@ 4.5. Economy-driven engagement
Pak’s projects deliberately intertwine economic mechanisms with artistic interaction, turning market dynamics into part of the artistic experience. In many cases, the engagement of the audience is driven by financial gameplay—bidding wars, token burns, price competition—such that market participation becomes a form of performance. The Sotheby’s sale The Fungible is a paradigmatic example: it deployed a complex auction structure with open editions, surprise price hikes, and contests for unique rewards, effectively a “gamified” economic performance. This not only generated approximately \$16.8 million in sales, but also created a frenzy in which “more than 3,000 individuals decided to join,” drawn in by the competitive game itself. The mechanics of scarcity and reward (e.g., top buyers receiving special NFTs) were designed to incentivize maximal participation. Here the market behavior was the art: once “the game and competition incentives disappear,” what remains are tokens whose value was defined by that performative context. Pak thus highlights, even cynically, the extent to which economic structures confer meaning in crypto art (Poposki, 2025).
@@
In Pak’s oeuvre, the act of buying or transacting becomes a critical ritual and medium. The Title (2021) consisted of identical images sold at different prices, making price itself the distinguishing feature of the artwork. One edition titled “Unsold” was listed for \$1 million and intentionally left unsold—a conceptual gesture underlining how value in art can be a product of hype and perception. This recalls Yves Klein’s 1957 exhibition of identical blue canvases sold at varying prices; as Klein observed, “the price…legitimately changes the experience of the work” despite no physical difference. Similarly, Pak has argued that an NFT is like a currency—judging it by the image alone is as foolish as picking a banknote. By merging art and economics, Pak invites collectors to reflect on their own motivations—are they driven by aesthetic appreciation or speculative gain? The economy around the artwork becomes integral to its narrative, effectively turning collectors into players in a financial drama. In Pak’s protocol art, market engagement is not a byproduct but a core feature that actively shapes the work itself.
@@ 4.6. Poetic message embedding in interaction rituals
Beyond the technical and economic layers, Pak’s works often carry metaphorical or poetic messages that are realized through the audience’s interactive rituals. In other words, the process a participant must follow is itself laden with symbolic meaning. Critic A.V. Marraccini notes that Pak treats the NFT medium as “a way to engage audiences in participatory conceptual art”—the concepts emerge through what the viewers or users are made to do. For instance, burn.art turns the destructive act of burning NFTs into a creative ritual. The platform’s mantra “Burn art to get ashes to get art to burn art” is a circular poem in itself, invoking themes of death and rebirth. Participants enact a cycle of sacrifice and renewal: by destroying one token, they generate another (the \$ASH token), which can then be spent on new art or even burned again. This cyclical interaction is a performative allegory about value and transformation—literally illustrating creation-through-destruction in a way that words alone could not.
Pak’s collaborations also embed messages in their structure. Censored (2022), created with Julian Assange, is explicitly about free expression and censorship. The project had two parts—one a dynamic single edition (Clock) counting Assange’s days behind bars, and the other an open invitation for people to speak. By telling users that Censored is “about you” and allowing them to tokenize any message they wish, Pak made the audience’s personal expression the core of the piece. The very name Censored reminds participants that their recorded messages might be obscured or blacked out, forcing reflection on the power of speech under observation. In Lost Poets, Pak evokes Jorge Luis Borges’s Library of Babel: an infinite library of all possible texts. The project “enlists participants to imagine” the lost contents of that library by naming their Poet NFTs and eventually inscribing words onto them. The ritual of naming and writing thus becomes a meditation on how meaning is generated—each user’s creative act is a fragment of a larger, hidden poem. Pak uses these game-like interactions not just as gimmicks but to convey ideas: the audience realizes the “poetic” concept by performing it. This approach recalls aspects of Fluxus or conceptual art, where simple actions (publishing a message, paying an exorbitant sum) become symbolic acts. The blockchain in Pak’s work provides a global, participatory stage for such rituals, ensuring that each interaction—each burn, each tokenized word—is both an artwork and a narrative gesture in Pak’s conceptual universe.
@@ 4.7. Interoperability enabling composability for emergence
Finally, Pak’s protocol-based approach leverages the interoperability of blockchain tokens to compose complex emergent ecosystems that transcend any single artwork. In blockchain terms, assets and smart contracts are composable: they can interact, integrate, or build upon each other permissionlessly. Pak exploits this composability to create an interlinked network of projects. The clearest example is the \$ASH token ecosystem. After The Fungible, Pak introduced \$ASH as a cryptocurrency earned by burning NFTs on burn.art. This token was not confined to one piece, but became a connective thread between works—a meta-artwork in its own right. Owners of \$ASH were rewarded in subsequent projects: before launching Lost Poets, Pak took a snapshot and airdropped 7,586 Pages to collectors holding at least 25 \$ASH, thus bridging the burn.art economy with the new literary NFT game. Conversely, once Lost Poets concluded, those Poet NFTs could be burned for \$ASH, feeding value back into the burn.art system. In this way, Pak created a feedback loop: one artwork’s output becomes another’s input. The emergent whole is an ecosystem of NFTs and tokens co-evolving across Pak’s oeuvre.
@@ 5. Discussion
5.1. Characteristics of Pak’s Protocol Artworks
@@
This study’s exploration of Pak’s protocol artworks reveals seven interlinked characteristics that distinguish protocol art as a practice. (C1) These works are fundamentally system-centric rather than object-centric. Their artistic meaning emerges from dynamic interconnectedness via protocols instead of static artifacts. (C2) As Galloway claimed “protocol is how control exists after decentralization” (Galloway, 2001), protocol arts employ autonomous governance for open-ended control, often via self-executing code or decentralized mechanisms, so that the artwork’s evolution is not predetermined by a single author but can continue to unfold indefinitely. (C3) This leads to distributed agency and communal authorship, as creative influence is spread across many participants and technical actants rather than residing in one creator, aligning with the idea that agency (and thus authorship) is shared among human and non-human actors in a network. (C4) Moreover, these artworks exhibit temporal dynamism and lifecycle aesthetics: they develop through time, with phases of growth or change designed into the experience, so that process and feedback become central to their aesthetic. (C5) Engagement with such works is often economic-driven, integrating incentive structures (for example, token rewards or market dynamics) that actively shape how audiences interact—a phenomenon noted by art scholars observing the immediate economic incentives at play in crypto art. (C6) At the same time, there is a deliberate embedding of poetic messages in interaction rituals: the very actions participants perform (trading, voting, combining elements, etc.) are imbued with symbolic meaning, turning user interactions into ritualized performances that carry the artwork’s conceptual message. (C7) Finally, protocol artworks leverage interoperability enabling composability for emergence, meaning they are designed to plug into larger ecosystems of code and community; like modular components, they can be reconfigured or linked with other protocols, allowing unexpected behaviors and creative outcomes to emerge from these combinations. Collectively, these seven characteristics illustrate how protocol art shifts the locus of art-making from singular objects to evolving systems of interaction, value and meaning.
@@
By situating these findings in broader intellectual contexts, we can appreciate protocol art as part of an interdisciplinary conceptual lineage. The system-centric focus strongly resonates with complex systems theory and the “systems aesthetics” tradition in art: as Jack Burnham observed, modern art increasingly “does not reside in material entities, but in relations between people and between people and the components of their environment”. In protocol art, the artwork is essentially a complex adaptive system (Buckley, 2017)—with inputs, outputs, and feedback loops—which accords with theoretical models where adaptive behavior and emergent order arise from networked interactions over time. Its open-ended governance and temporal evolution connect to the field of artificial life, which seeks systems that continually produce novelty rather than reaching a fixed end state. This open-endedness (Stanley et al., 2017), coupled with deliberate feedback mechanisms, echoes the temporal dynamics of complex adaptive systems studied in complexity science, reinforcing how unpredictable, ongoing change becomes an aesthetic virtue. Meanwhile, the ethos of distributed agency and communal authorship in these works finds precedent in media theory and anthropology: for instance, actor-network theory contends that creative agency is distributed across a web of human and non-human actants, not concentrated in an autonomous individual. Similarly, in interactive and participatory art literatures, authorship is often viewed as an emergent, collective process—the community of participants effectively co-creates and continually reshapes the piece (Bishop, 2023). The inclusion of economic-driven engagement situates protocol art in dialogue with social computing and crypto-economic systems: just as Web3 platforms rely on token-based incentives to drive user participation and loyalty, protocol artworks harness economic game dynamics as artistic material, blurring the line between aesthetic experience and market behavior. This integration of economic and social incentives also invokes evolutionary game theory within complex systems, where value-based choices influence the system’s trajectory. Furthermore, the poetic messaging through interaction rituals can be interpreted via performance studies and ritual theory in media. Even routine or rule-bound interactions carry expressive and symbolic weight—indeed, ritualized acts are “anything but purposeless” and can “constitute social reality” through shared symbolic action. In protocol art, the choreography of user interactions (such as collective decision-making or repetitive transactions) functions as a form of narrative or commentary, akin to a ceremonial performance that conveys meaning beyond its practical function. Finally, the principle of interoperability and composability aligns with discussions in software and platform studies about modular design and emergent behavior. In blockchain-based art, for example, smart contracts are conceived as interoperable building blocks, combinable like Lego pieces to create novel structures (Sun et al., 2025). This composability not only reflects a technical design philosophy but also fosters creative emergence: new artistic forms and communities can arise by linking protocols together, much as complex behaviors emerge when simple units interact in networked ecologies. Linking these characteristics to established frameworks thus shows that protocol art is not an isolated novelty but converges with long-running threads in complexity science, new media art, and socio-technical theory, reframing them in a unique artistic context.
@@ 5.2.1. Beyond Conceptual Art
“The idea becomes the machine that makes the art,” wrote Sol LeWitt in 1967 (LeWitt, 1967)—a formulation that conceptual art could only fulfill metaphorically, since LeWitt’s wall drawings still required human execution and institutional presentation. Pak’s protocol art literalizes this aspiration: the smart contract is the machine, and the idea—encoded once as immutable bytecode—executes itself without further authorial intervention. Joseph Kosuth’s One and Three Chairs (1965) established that an idea can eclipse the physical object when reframed through language and documentation. Pak embraces the primacy of ideas but transports them into a rule-executing substrate. The Title (2021), whose only variables are price and on-chain title, echoes Kosuth’s linguistic gambit, yet its value is continuously arbitrated by decentralized markets rather than by academic discourse or institutional framing. Because the smart contract is immutable, its “statement” cannot be re-contextualized by curators; it is enacted in perpetuity by the network. Thus Pak’s work is not merely an illustration of a concept but a permanently operational logic that strangers must literally transact with. The artwork lives as executable code—a self-verifying proposition—rather than a documented proposition that awaits institutional re-presentation.
@@ 5.2.3. Beyond Participatory / Crowdsourcing Art
Relational-aesthetic projects of the 1990s (Bourriaud) invited viewers to co-produce meaning through convivial encounters, but those encounters were typically ephemeral and locally bounded. Pak’s participation is ledger-bound. In Lost Poets (2021–22), each name or verse a collector inscribes becomes an indelible block in a public archive—a form of distributed authorship impossible in earlier participatory works whose traces remained in wall labels or catalogs. Crucially, participation is also economically weighted: burning a Page to earn words entails an explicit cost, so creative decisions carry financial risk. This introduces game-theoretic depth absent from classic crowdsourcing art (e.g., MTurk drawings or open Wikis), aligning Pak’s practice with emerging literatures on cryptoeconomic design rather than with social-practice documentation alone. Moreover, Censored (2022) demonstrates that protocol-based participation can transcend the art world entirely. AssangeDAO—a collective of over 10,000 members—pooled 16,593 ETH (approximately \$52.8 million) to acquire the Clock NFT, with proceeds directed to Julian Assange’s legal defense via the Wau Holland Foundation. When Assange was released in June 2024, the participation had demonstrably contributed to a real-world political outcome—a reach no gallery-bound participatory artwork has achieved.
@@ 5.2.4. Beyond Interactive Art
Interactive art from Myron Krueger’s Videoplace to Rafael Lozano-Hemmer’s public-space works invites real-time feedback loops, but those loops terminate once the exhibition ends. Protocol art designs interaction into on-chain fact. When a collector merges two Merge tokens or burns an NFT for \$ASH, the blockchain records a state change that affects every future viewer and holder. Interaction ceases to be a momentary spectacle and becomes infrastructure—a permanent mutation in the artwork’s ontology. Each transaction—a merge, a burn, a tokenized message in Censored—is an irreversible on-chain fact, not a fleeting gesture that vanishes when the gallery closes. Moreover, protocol interaction is asynchronous and scalable: millions of transactions are handled by the Ethereum network itself, so participation is bounded only by the scale of the chain rather than by gallery capacity or curatorial gatekeeping.
@@ 5.2.5. Beyond Performance Art
Performance art often foregrounds the performative encounter between artist and participant—think of Yoko Ono’s Cut Piece (1964) (Ono, 1964), in which audience members were invited to cut away the artist’s clothing, making collective action the medium. Protocol art extends this logic beyond the gallery and across the entire internet: anyone, anywhere, can permissionlessly participate in the work by signing a transaction, burning a token, or inscribing a message—no invitation, no curator, no physical co-presence required. Where Ono’s performance was bounded by a single room and a single sitting, Censored (2022) enrolled nearly 30,000 participants across two days and multiple continents, each contributing a locked message that became part of an enduring on-chain chorus. The protocol replaces the artist’s body as the performing agent: it executes its rules indefinitely, delegating durational endurance to automated smart-contract logic and DAO stewardship rather than to flesh and presence. Thus the locus of performance shifts from a singular body in time to a permissionless, planetary-scale system in perpetuity.
5.3. Limitations and Future Work
Our intensive case study of Pak’s nine landmark protocol artworks demonstrates how smart-contract logic can serve as an artistic medium while providing foundational vocabulary for the emerging field. By focusing on the paradigmatic creator of protocol art, we have articulated seven characteristics that can now serve as analytical lenses for future comparative research.
This study does not claim that all protocol art exhibits the seven characteristics we identify; such a claim would require cross-artist sampling inappropriate to our exploratory aims. Instead, we offer a generative contribution: a vocabulary and analytical framework derived from close engagement with protocol art’s most influential practitioner. These characteristics function as what Blumer termed sensitizing concepts—interpretive devices that orient attention without predetermining findings—that future researchers can apply, adapt, or reject when analyzing other artists (Blumer, 1954). The value of our contribution lies not in statistical generalization but in conceptual articulation: naming and describing phenomena that lacked prior theorization. Whether terra0, Sarah Friend, or Rhea Myers exhibit similar characteristics is an empirical question our framework now makes askable. Skeptics have questioned whether NFTs substantially modify art-world dynamics at all (Radermecker and Ginsburgh, 2023), making such comparative work all the more pressing. We anticipate that some characteristics (e.g., system-centric composition, autonomous governance) will prove broadly applicable to protocol art as a genre, while others (e.g., economy-driven engagement at Pak’s scale) may be specific to commercially-oriented practice. Comparative research is needed to distinguish genre-defining features from artist-specific idiosyncrasies.
At the same time, this focus exposes the field’s present fragmentation: there is still no shared vocabulary, taxonomy, or methodological toolkit robust enough to account for the full diversity of protocol-driven practice. We therefore call for a concerted, interdisciplinary effort to systematize knowledge around protocol art. Concretely, future work should